Will someone please help me with this probelm. Im trying to solve for x:
Log Base x of 8/27=-3. Can someone help with the answer. And also please show me how to do it.

Answer by stanbon(48535)

Take the anti-log to get x^(-3)=8/27
Raise both sides to the -1/3 power to get:
x=(8/27)^(-1/3)
The negative in the exponent means to invert, so:
x=(27/8)^(1/3)
The 1/3 means to take the cube root, so:
x=3/2
Cheers,
Stan H.